Arve ramp-up

Lassi Tuura proposed the “ramp-up” procedure at the Edinburgh workshop (see the minutes, with hyperlink to his slides). The principle was accepted.

Also at Edinburgh, Arve was proposed as a candidate for ramp-up, with the first point, qualification already accepted, leaving the second point, timescale, for evaluation.

Three evaluators (dell’Acqua, Candlin, Tuura) were appointed and reported at the end of June. LAT was the first, with a full report, summarized on the next slide.

Ad’A and DJC largely agreed with him, though DJC differed on one important point (following slide).

• General: apply style and other guidelines • Decomposition: too much interdomain vis-

ibility, singletons, inheritance

• Graphics: too much mixed with GUI (use a canvas)

• Event gen/sim:

– leave Gismo in (untouched) until GEANT4 – Design of digi storage mechanism

• Control/GUI:

– embed GUI in control, instead of control in GUI

– replace control with another design Time estimate for all this: end March 1998 DJC made these comments or variations:

• General: apply style and other guidelines *** this will not only change Arve, but will probably react back on the guidelines.

– Design of digi storage mechanism *** this is a development, not within but after rampup

• Control/GUI:

– embed GUI in control, instead of control in GUI *** why cannot they be at equal level?

– replace control with another design *** definitely not, within the rampup; this is a major enhancement, afterwards

Time estimate for all this: end September 1997